# references.bib

### v3

@@ 6.1 Pro- and Ind-Completions of categories of finite objects
For categories of algebras, the Pro-completion can equivalently be seen as consisting of the profinite algebras of the given type, whilst the Ind-completion can be seen as obtaining all algebras of that same type. As such, we have that:
@@
where $\mathbf{ImFinPos}$ is the category of image-finite posets with p-morphisms. And indeed, we have that such a category is, by the main results from [Bezhanishvili2008], dual to the category of profinite Heyting algebras, as desired. Indeed, a similar picture to this one has been discussed for finite modal algebras, and their duals, finite Kripke frames, in the recent preprint [profinitenessmonadicityuniversalmodels], where the above dualities are briefly sketched.
@@ 6.2 An Adjunction between the category of Image-Finite posets and Posets with Monotone Maps
We now prove the following, which is analogous to the property from Lemma 18:
Lemma 57.
@@
Using this we can show that the poset $P_{G}(X)$ will be free in the desired way. The key fact that is missing concerns the fact that, after applying Lemma 57 infinitely often, one needs, as in Proposition 20, to show that the lifting is indeed a p-morphism. There we used compactness, which certainly does not hold for arbitrary posets.
Proposition 58.
@@
Using similar arguments as those of Theorem 22, we can then show:
Proposition 59.
@@
It is also notable that the above methods cannot be expected to be extended to yield an adjunction between $\mathbf{Pos}_{p}$ and $\mathbf{Pos}$ as above. Indeed, as noted in [openmapsdoesnothaveproducts], $\mathbf{Pos}_{p}$ does not have binary products. If $I:\mathbf{Pos}_{p}\to\mathbf{Pos}$ had a right adjoint, then all limits in $\mathbf{Pos}$ would have to be preserved, which is absurd.
@@ 7 Conclusions and Further Research
As noted throughout the text, the material presented here does not exhaust the range of applications of these constructions and the ideas wherein: in [openmapsandinterpolantcomplexity] we expect to present proofs of uniform interpolation, as well as new results concerning interpolant complexity, and the structure of free Heyting algebras, exploiting the description of coproducts given above. Moreover, this research leaves open several questions, both of a technical and of a conceptual nature.
One fact which we have not worked with concerns the structure of the Priestley spaces obtained by applying $V_{g}(-)$; when $X$ i a finite poset, this yields a finite poset, which has a natural structure as a (bi-)Esakia space. It is not clear that even starting with a bi-Esakia space $X$, $V_{g}(X)$ will be bi-Esakia; we have been able to show that $V_{r}(X)$ will be co-Esakia, but all other facts remain open.

### v4

@@ 7.1 Pro- and Ind-Completions of categories of finite objects
For categories of algebras, the Pro-completion can equivalently be seen as consisting of the profinite algebras of the given type, whilst the Ind-completion can be seen as obtaining all the locally finite algebras of that same type – which in the case of Boolean algebras or Distributive lattices, will be all algebras. As such, we have that:
@@
where $\mathbf{ImFinPos}$ is the category of image-finite posets with p-morphisms; this is, indeed, the main result from [Bezhanishvili2008], where it is shown that these are dual to the category of profinite Heyting algebras, as desired. Indeed, a similar picture to this one has been discussed for finite modal algebras, and their duals, finite Kripke frames, in the recent work [profinitenessmonadicityuniversalmodels], where the above dualities are briefly sketched.
@@ 7.2 An Adjunction between the category of Image-Finite posets and Posets with Monotone Maps
We now prove the following, which is analogous to the property from Lemma 3.9:
Lemma 7.2.
@@
Using this we can show that the poset $P_{G}(X)$ will be free in the desired way. The key fact that is missing concerns the fact that, after applying Lemma 7.2 infinitely often, one needs, as in Proposition 3.11, to show that the lifting is indeed a p-morphism. There we used compactness, which certainly does not hold for arbitrary posets, and this is the place where image-finiteness becomes crucial.
Proposition 7.3.
@@
We can then show:
Proposition 7.4.
@@
Proof.
One verifies, analogously to Lemma 3.12 that the construction $P_{G}^{g}$ is functorial, and the rest follows just like in Theorem 3.13. ∎
@@
It is also notable that the above methods cannot be expected to be extended to yield an adjunction between $\mathbf{Pos}_{p}$ – the category of posets with p-morphisms – and $\mathbf{Pos}$ as above. Indeed, as noted in [openmapsdoesnothaveproducts], $\mathbf{Pos}_{p}$ does not have binary products; on the other hand, if $I:\mathbf{Pos}_{p}\to\mathbf{Pos}$ had a right adjoint, then all limits in $\mathbf{Pos}$ would have to be preserved.
We conclude by noting that the two constructions provided – $V_{G}$ and $P_{G}$ – are indeed intimately related for any finite poset $P$. For this we will need a few more technical developments:
Definition 7.5.
Let $(V_{n}(X))_{n\in\omega}$ be a Vietoris complex. Let $x\in V_{n}(X)$ be any point. We say that $x$ is prestable if $r_{n+1}^{-1}[x]$ is a singleton; we say that it is stable if whenever $x\leq x^{\prime}$, then $x^{\prime}$ is prestable [^8] These concepts seem to have been first considered by Dito Pataraia in unpublished work. I credit Pataraia for the main ideas of these results, though the proofs are slightly different..
Lemma 7.6.
Assume that $x\in X_{n+1}$. If $r_{n+1}(x)$ is stable, then $x$ is stable as well.
Proof.
Assume that $C,D$ are such that $r_{n+2}(C)=r_{n+2}(D)=x$. Suppose that $y\in C$. Then $x\leq y$. Since $D$ is $r_{n+1}$-open, there is some $y^{\prime}\in D$ such that $r_{n+1}(y)=r_{n+1}(y^{\prime})$. Since $r_{n+1}(x)$ is stable, then $y=y^{\prime}$, so $y\in D$. This shows that $C=D$, which, since ${\uparrow}x$ roots at $x$, proves that $x$ is prestable. The same argument shows that all successors of $x$ are prestable, obtaining the result. ∎
Lemma 7.7.
[Pataraia’s Stability Lemma] Let $x\in V_{n}(X)$ be arbitrary. Then $x^{\bullet}=\{{\uparrow}({\uparrow}y):x\leq y\}\in X_{n+2}$ is stable, and $r_{n+1}(r_{n+2}(x))=x$.
Proof.
Note that $x^{\bullet}$ is $r_{n+1}$-open. Now assume that $r_{n+3}(C)=r_{n+3}(D)=x^{\bullet}$. Let $Y\in C$; then $x^{\bullet}\leq Y$. This means that $r_{n+2}(Y)={\uparrow}({\uparrow}y)$ for $x\leq y$. Now suppose that $y\leq z$. Then note that ${\uparrow}({\uparrow}z)\geq r_{n+2}(Y)$; since the latter is $r_{n+1}$-open, there is some $K\in Y$ such that $r_{n+1}(K)={\uparrow}z$. Note that $Y^{\prime}=Y\cap{\uparrow}K$ is then such that $x^{\bullet}\leq Y^{\prime}$, so by the same argument, $K={\uparrow}({\uparrow}m)$ for some $m$; but then given its root, $K={\uparrow}({\uparrow}z)$. Hence $Y=y^{\bullet}$.
Note that by $r_{n+2}$-openness of $D$, there is some $Y^{\prime}\in D$ such that $r_{n+2}(Y^{\prime})=r_{n+2}(Y)$. The exact same argument then shows that $Y^{\prime}=y^{\bullet}$. So $Y\in D$, which shows that $C=D$. This shows that $x^{\bullet}$ is prestable. Now note that then, by the argument we have just shown, any successor of $x^{\bullet}$ is of the form $y^{\bullet}$, so the former is stable. ∎
Corollary 7.8.
If $n\in\omega$ is arbitrary, and $x\in V_{n}(X)$, there is some point $x^{\prime}\in V_{G}(X)$ such that ${\uparrow}x^{\prime}$ is finite, and $x^{\prime}(n)=x$.
Proof.
Using Lemma 7.7 and Lemma 7.6, we can construct an extension of $x$ to a point of $V_{G}(X)$ which only has finitely many extensions, using the fact that the root map is an isomorphism on the restriction from $x^{\bullet}$ to $x$, and onwards along the construction. ∎
Theorem 7.9.
Let $P$ be a finite poset. Then the image finite part of $V_{G}(P)$ is dense in the space.
Proof.
For this we need to show that for each $U\subseteq V_{n}(P)$ a (clopen) set, $\pi_{n}^{-1}[U]$ contains a point $x$ which only sees finitely many points. Indeed, suppose that $x(n)\in U$. Using Corollary 7.8, let $\overline{x}$ be a finite point such that $\overline{x}(n)=x(n)$. Then $\overline{x}\in\pi_{n}^{-1}[U]$. This shows that the image-finite part is dense, as desired. ∎
Recall that given a poset $(P,\leq)$, an ordered-topological space $(X,\leq,\tau)$ and an order-preserving map $p:P\to X$, we say that $p$ is an order-compactification if $p$ is an order-homeomorphism onto the image, and $p[P]$ is dense in $X$ (see [nachbin1965topology]). Then we have:
Corollary 7.10.
For each finite poset $P$, $V_{G}(P)$ is an order-compactificationof $P_{G}(P)$.
Proof.
By Theorem 7.9 this follows immediately. ∎
@@ 8 Conclusions and Further Research
As noted throughout the text, the material presented here does not exhaust the range of applications of these constructions and the ideas wherein: in ongoing work [AlmeidaBezhanishviliDukic], the results of the present paper have applications in the study of coalgebraic logic; in future work we also expect to be able to use these ideas to investigate the structure of normal forms in superintuitionistic logics, as well as issues such as the existence and complexity of interpolants. Moreover, the results of this paper leave open several questions, both of a technical and of a conceptual nature.
One fact which we have not worked with concerns the structure of the Priestley spaces obtained by applying $V_{g}(-)$; when $X$ i a finite poset, this yields a finite poset, which has a natural structure as a (bi-)Esakia space; when $X$ is linear, then $V_{r}(X)$ is homeomorphic to $V(X)$, and hence by the results of Section 5, is an Esakia space. But it is not clear that even starting with a bi-Esakia space $X$, $V_{g}(X)$ will be bi-Esakia.